OpenStudy (anonymous):

consider a 67-g chunk of ice (H = 6.0 kJ/mol) in a beaker immersed in a water bath. To produce just enough heat to melt the ice, how many moles of solid NaOH (H = –445.1 kJ/mol) must you dissolve in the water bath?

OpenStudy (chmvijay):

40 gram liberates -445 KJ 18 gram of water(ice) needs around 6 Kj of energy but you have 37 gram of ice chunk =6*37 /18 = total 12.33 KJ is needed to liberate -445 Kj you need 40 gram of NaOH to liberate -12.33 KJ you need =40*12.33 /445 =1.10 gram of NaOH